An object has an applied force of 323 N upwards acting on it and accelerates downwards at 6.0 m/s2. What is the mass of this object?
1
Expert's answer
2019-05-24T09:57:06-0400
"ma=mg-F"

"m=\frac{F}{g-a}=\frac{323}{9.8-6}=85\ kg"